# Segmented Operations using Matrix Multiplications

### v1

@@ A.1 Matrix multiplication circuits
A.3 Reverting speculative scans using (masked) vector operations
@@
First, Algorithm 1 is simple and provides insight about the vector computations required. Its main drawback is that it performs reductions of $\mathcal{O}(s)$ numbers for each segment, hence requiring $\mathcal{O}(\log(s))$ steps and $\mathcal{O}(ns)$ work. To further reduce the work overhead of the speculation, we present Algorithm 2, which has $\mathcal{O}(n\log(s))$ work and $\log(s)+\mathcal{O}(1)$ steps. The key difference of the two algorithms is that the latter uses a Kogge–Stone prefix masked reduction to avoid the work overhead (recomputation) compared to Algorithm 1.
@@
5:         if $\mu==0$ or ($\mu==1$ and $\overline{\bm{\mathrm{f}}}_{s}(0)==1$then
@@
8:              $\bm{\mathrm{z}}_{s}\leftarrow\bm{\mathrm{x}}_{s}$triangleright $s$ segments, no scan required. Revert.
@@
1:procedure RevertSpecMMVRAM($\bm{\mathrm{z}}$, $\bm{\mathrm{x}}$, $\bm{\mathrm{f}}$, $s$) $\triangleright$ Works in-place on $\bm{\mathrm{z}}$
@@
5:         if $\mu==0$ or ($\mu==1$ and $\overline{\bm{\mathrm{f}}}_{s}(0)==1$then
@@
8:              $\bm{\mathrm{z}}_{s}\leftarrow\bm{\mathrm{x}}_{s}$triangleright $s$ segments, return input.
@@
12:         $\bm{\mathrm{w}}_{s}\leftarrow\mathsf{SHIFTR}\left(\mathsf{MASK}\left(\bm{\mathrm{z}}_{s},\bm{\mathrm{q}}_{s}\right),1\right)$triangleright Correction vector.
@@
A.4 Vector differentiation using matrix multiplications
Here, we demonstrate that it is possible to use matrix multiplication to compute vector differentiation in the last step of Algorithm 1. Let $\bm{\mathrm{A}}$ be a row-wise matrix view with $s$ columns of an arbitrary vector $\bm{\mathrm{x}}$ (padded accordingly). Let
$\bm{\mathrm{D}}_{s}:=\begin{bmatrix}1&-1&0&\dots&0\\ 0&1&-1&\dots&0\\ 0&0&1&\dots&0\\ \vdots&0&0&\ddots&-1&\\ 0&0&\dots&0&1\end{bmatrix}_{s\times s}.$ (4)
It is easy to verify that the matrix product $\bm{\mathrm{A}}\bm{\mathrm{D}}_{s}$ computes the difference $x_{i}-x_{i-1}$ on each $s$-block except the boundary entries between blocks. The remaining entries on the boundary can be computed using the VCU.
A.5 Dataset details
Table A.1 shows the matrix characteristics, including matrix sizes and sparsity statistics of the matrices used in Section 7.
Table A.1: Sparse Suite Matrices.
Name Dimensions $nnz$ $nnz$ per row in $nnz$ per row avg $nnz$ per row max
Protein 36K $\times$ 36K 4.3M 2K 2K 2K
Epidemiology 526K $\times$ 526K 2.1M 18 119 204
Economics 207K $\times$ 207K 1.3M 1 6 44
Circuit5M 5.6M $\times$ 5.6M 2.1M 18 119 204
ASIC_680K 683K $\times$ 683K 3.9M 1 6 395K

### v2

@@ A.2 Block-Recursive Segmented Scan Algorithm
To use Lemma˜A.1 in Algorithm 1, recall that the speculation prepares the entire vectors $\overline{\bm{\mathrm{x}}}\leftarrow\mathsf{MATMUL}(\bm{\mathrm{x}},\bm{\mathrm{U}}_{s})$ and $\overline{\bm{\mathrm{f}}}\leftarrow\mathsf{MATMUL}(\bm{\mathrm{f}},\bm{\mathrm{U}}_{s})$, which contain $\mathcal{O}(n/s)$ blocks of size $s$. By concatenating $\lceil n/s\rceil$ copies of the circuit of Lemma A.1, we can build a uniform family of circuits that implements the desired vector instruction $\mathsf{REVSPEC}$. The total circuit size is $\mathcal{O}\left(\frac{n}{s}(s^{2}B+B^{2})\right)=\mathcal{O}\left(n(sB+B^{2}/s)\right)$, and the depth remains constant after concatenation.
In the last phase of Algorithm 1, following the recursive call, the VCU are invoked to scalar-broadcast-vector-add the previous block’s last entry (scalar) to the current block’s first segment (UpdateFirstSegment in Algorithm˜1).
Figure˜A.1 depicts an example of an execution of Algorithm˜1 for an input of size $16$ and block size $s=4$. In this example, matrix multiplication computes the unsegmented scans, and the misspeculated entries are coloured red. The reversion of the speculation is followed before the recursions take place. In the final step, vector-masked operations are employed to propagate the last entry of each block to the first segment of the next block.
Remark A.2.
The circuit of Lemma A.1 is ideal for our theoretical analysis, but it is rather complicated compared to the rest of the vector instructions in Table˜A.2. In a realistic setting, it would need to be implemented as a dedicated circuit inside the VCU, which can be prohibitive in terms of area. In Section˜A.2.1 we describe an alternative, Algorithm˜3, which simulates the $\mathsf{REVSPEC}$ instruction using only a few standard vector instructions, which are typically implemented in existing architectures.
A.2.1 Reverting speculative scans using (masked) vector operations
@@
First, Algorithm˜2 is simple and provides insight about the vector computations required. Its main drawback is that it performs reductions of $\mathcal{O}(s)$ numbers for each segment, hence requiring $\mathcal{O}(\log(s))$ steps and $\mathcal{O}(ns)$ work. To further reduce the work overhead of the speculation, we present Algorithm˜3, which has $\mathcal{O}(n\log(s))$ work and $\log(s)+\mathcal{O}(1)$ steps. The key difference of the two algorithms is that the latter uses a Kogge–Stone prefix masked reduction to avoid the work overhead (recomputation) compared to Algorithm 2.
@@
5:    if $\mu==0$ or ($\mu==1$ and $\overline{\bm{\mathrm{f}}}_{s}(0)==1$) then
@@
8:     $\bm{\mathrm{z}}_{s}\leftarrow\bm{\mathrm{x}}_{s}$ $\triangleright$ $s$ segments. Revert.
@@
1:procedure RevertSpecVector($\bm{\mathrm{z}}$, $\bm{\mathrm{x}}$, $\bm{\mathrm{f}}$, $s$) $\triangleright$ Works in-place on $\bm{\mathrm{z}}$
@@
5:    if $\mu==0$ or ($\mu==1$ and $\overline{\bm{\mathrm{f}}}_{s}(0)==1$) then
@@
8:     $\bm{\mathrm{z}}_{s}\leftarrow\bm{\mathrm{x}}_{s}$ $\triangleright$ $s$ segments. Revert.
@@
12:    $\bm{\mathrm{w}}_{s}\leftarrow\mathsf{SHIFTR}\left(\mathsf{MASK}\left(\bm{\mathrm{z}}_{s},\bm{\mathrm{q}}_{s}\right),1\right)$ $\triangleright$ Correction vector.
@@
A.2.2 VCU instructions
Table˜A.2 summarizes the vector instructions used for Algorithm 1, which achieves the best work for segmented scan (reported in Theorem 4.3). All these instructions can be implemented with uniform AC[0] circuits for every input size $n$ as follows.
  • Element-wise $\mathsf{ADD}/\mathsf{SUB}$ can also be achieved with AC[0] circuits, with $\mathcal{O}(nB^{2})$ and $\mathcal{O}(B)$ fan-in boolean gates. See e.g. [61, Chapter 1] or [21, Chapter 8] for further details and for more advanced circuits and techniques. We note that, in the main algorithm, we only use subtractions aba-b for the special case where $a\geq b$.
  • $\mathsf{SCATTERS}/\mathsf{GATHERS}$ are also parametrized by $s$. As the name suggests, $\mathsf{GATHERS}$ copies the last element of the $i$-th $s$-segment of $\bm{\mathrm{x}}$, to $\bm{\mathrm{z}}(i-1)$. Accordingly, $\mathsf{SCATTERS}$ copies the (i1)(i-1)-th element of $\bm{\mathrm{x}}$ to the last position of the $i$-th $s$-segment of $\bm{\mathrm{z}}$. The corresponding circuits have size $\mathcal{O}(sB)$, since they only copy one element per segment to a single position in the output, and the fan-in/-out of the gates are all constant.
  • Finally, the instruction $\mathsf{REVSPEC}$ is a uniform circuit that corrects the misspeculated (unsegmented) scans in Algorithm 1. This circuit is described in detail in Section 4, specifically, in Lemma A.1.
Table A.2: Circuit sizes and fan-in for vector operations of length $n$ used in the main Algorithm˜1. Integers (int) have bit-width $B$. All circuits have constant depth, and $s$ is the fixed model parameter. All instructions (except $\mathsf{REVSPEC}$) are either present or easy-to-implement in existing architectures.
Operation Size Fan-in / -out Description
1. $\bm{\mathrm{z}}:\text{vec[bit]}\leftarrow\mathsf{AND}/\mathsf{OR}(\bm{\mathrm{x}}:\text{vec[bit]},\bm{\mathrm{y}}:\text{vec[bit]})$ $\mathcal{O}(n)$ $\mathcal{O}(1)$ $\bm{\mathrm{z}}(i)\leftarrow\bm{\mathrm{x}}(i)\wedge\bm{\mathrm{y}}(i)$ (resp. $\bm{\mathrm{x}}(i)\vee\bm{\mathrm{y}}(i)$).
2. $\bm{\mathrm{z}}:\text{vec[bit]}\leftarrow\mathsf{NOT}(\bm{\mathrm{x}}:\text{vec[bit]})$ $\mathcal{O}(n)$ $\mathcal{O}(1)$ $\bm{\mathrm{z}}(i)\leftarrow\neg\bm{\mathrm{x}}(i)$.
3. $\bm{\mathrm{z}}:\text{vec[int]}\leftarrow\mathsf{MASK}(\bm{\mathrm{x}}:\text{vec[int]},\bm{\mathrm{y}}:\text{vec[bit]})$ $\mathcal{O}(nB)$ $\mathcal{O}(1)$ / $\mathcal{O}(B)$ $\bm{\mathrm{z}}(i)\leftarrow\bm{\mathrm{x}}(i)$ if $\bm{\mathrm{y}}(i)=1$, else $\bm{\mathrm{z}}(i)\leftarrow 0$.
4. $\bm{\mathrm{z}}:\text{vec[bit]}\leftarrow\mathsf{ISZERO}(\bm{\mathrm{x}}:\text{vec[int]})$ $\mathcal{O}(nB)$ $\mathcal{O}(B)$ / $\mathcal{O}(1)$ $\bm{\mathrm{z}}(i)\leftarrow 1$ if $\bm{\mathrm{x}}(i)=0$, else $\bm{\mathrm{z}}(i)\leftarrow 0$.
5. $\bm{\mathrm{z}}:\text{vec[int]}\leftarrow\mathsf{ADD}/\mathsf{SUB}(\bm{\mathrm{x}}:\text{vec[int]},\bm{\mathrm{y}}:\text{vec[int]})$ $\mathcal{O}(nB^{2})$ $\mathcal{O}(B)$ $\bm{\mathrm{z}}(i)\leftarrow\bm{\mathrm{x}}(i)+\bm{\mathrm{x}}(j)$ (resp. $\bm{\mathrm{x}}(i)-\bm{\mathrm{x}}(j)$).
9. $\bm{\mathrm{z}}:\text{vec[int]}\leftarrow\mathsf{REVSPEC}(\overline{\bm{\mathrm{x}}}:\text{vec[int]},\overline{\bm{\mathrm{f}}}:\text{vec[int] };s)$ $\mathcal{O}\left(nB(s^{2}+\frac{B}{s})\right)$ $\mathcal{O}(s+B)$ / $\mathcal{O}(sB)$ Reverts misspeculation (Lemma A.1).
A.3 Additional implementation details
A.3.1 Details on AscendC development
The programming model is asynchronous and pipeline-based, as multiple AI-core executions can be pipelined while data flows from one unit to the other, optimizing performance. Computational kernels could be written in AscendC, a programming language built on top of C++ allowing for fine-grained control over the hardware components of Ascend. As the accelerator specifically targets AI workloads, in which we can usually apply quantization to handle data sizes, it mainly offers support for int8, float16, and float32, making the accelerator suitable for both training and inference steps; see [38] for further details.
A.3.2 Implementation of Ascend operators for segmented scan and sum
A.3.3 SpMV implementation
Here, we provide more details about the SpMV algorithm. In Algorithm 4, we list the variant of Algorithm 4 that is implemented and used in our experimental evaluation and, most importantly, takes into account the corner case where a sparse matrix has empty rows; see [40, 41].
[…]
2:  $\bm{\mathrm{z}}\leftarrow$ CsrGather($\bm{\mathrm{A}}.\bm{\mathrm{val}}$, $\bm{\mathrm{A}}.\bm{\mathrm{col}}$, $\bm{\mathrm{x}}$)
3:  $\hat{\bm{\mathrm{z}}}\leftarrow$ Scan($\bm{\mathrm{z}}$)
4:  $\bm{\mathrm{w}}\leftarrow$ GatherSpMV($\hat{\bm{\mathrm{z}}},\bm{\mathrm{A}}.\bm{\mathrm{row}}$)
5:  return Diff($\bm{\mathrm{w}}$)
Algorithm 4 MMV-RAM Sparse CSR Matrix-Vector (SpMV)
Algorithm 5 describes the CsrGather step of our SpMV algorithm. Given $\bm{\mathrm{A}}$ and $\bm{\mathrm{x}}$ the algorithm compute $\bm{\mathrm{z}}$ such that:
$\bm{\mathrm{z}}(i):=\bm{\mathrm{val}}(i)*\bm{\mathrm{x}}(\bm{\mathrm{col}}(i))\quad 0\leq i<\operatorname{nnz}.$ (4)
In our implementation, the algorithm collects the whole vector $\bm{\mathrm{x}}$ into the scratchpad memory of each AIVs and then performs a gather and an element-wise multiplication. A current limitation of the algorithm is that it only works for input vectors $\bm{\mathrm{x}}$ of length up to 70K since the whole vector $\bm{\mathrm{x}}$ must be collected on each AIV core on the gather step (Line 4). Finally, the GatherSpMV Algorithm 6 implements a specialized step to collect the last element of each row after the SCAN step and add zeros in correspondence of the initial matrix empty rows.
1:procedure CsrGather($\bm{\mathrm{val}}$, $\bm{\mathrm{col}}$, $\bm{\mathrm{x}}$)
2:  $\bm{\mathrm{z}}\leftarrow$ Zeros vector of length $\operatorname{nnz}$
3:  for each $s$-tile ($\bm{\mathrm{val}}_{s},\bm{\mathrm{col}}_{s},\bm{\mathrm{z}}_{s}$) do
[…]
5:    $\bm{\mathrm{z}}_{s}\leftarrow$ Mul($\bm{\mathrm{val}}_{s}$, $\bm{\mathrm{w}}_{s}$) $\triangleright$ Element-wise.
6:  return $\bm{\mathrm{z}}$
Algorithm 5 CSR Gather step of [9], see also [40].
1:procedure GatherSpMV($\bm{\mathrm{y}}$, $\bm{\mathrm{row}}$) $\triangleright$ $\operatorname{\mathrm{len}}(\bm{\mathrm{y}})=\operatorname{nnz}$
2:  $s\leftarrow$ tiling parameter based on $\operatorname{nnz}$ and $m$.
3:  $(\bm{\mathrm{z}},\operatorname{nnz})\leftarrow$ HandleLeadingZeros($\bm{\mathrm{y}}$, $\bm{\mathrm{row}}$)
4:  $\bm{\mathrm{z}}(\operatorname{nnz}:)\leftarrow\text{GATHER($\bm{\mathrm{y}}$, $\bm{\mathrm{row}}$, $s$, offset = 1)}$
5:  return $\bm{\mathrm{z}}$
6:procedure HandleLeadingZeros($\bm{\mathrm{y}}$, $\bm{\mathrm{row}}$)
7:  Let $\bm{\mathrm{z}}\leftarrow\bm{\mathrm{y}}$, $i\leftarrow 0$
8:  while $\bm{\mathrm{row}}(i)==0$ do
9:   $\bm{\mathrm{z}}(i)=0$
10:   $i\leftarrow i+1$
11:  return $\bm{\mathrm{z}},i$ $\triangleright$ $i$: # of leading zero rows of $\bm{\mathrm{A}}$.
Algorithm 6 Specialized gather for SpMV.